In which of the following species is the oxidation number of sulfur less than 6?

A. SO3

B. Na2SO4

C. SO2


1
Expert's answer
2021-07-26T06:42:56-0400

A.

O: "-2 \times 3 = -6"

S: +6


B.

Na: "+1 \times 2 = +2"

O: "-2 \times 4 = -8"

S: +6


C.

O: "-2 \times 2 = -4"

S: +4

Answer: option C. SO2